# Lensing and wave optics in the strong field of a black hole

### v2

@@ Appendix A Convergence of the multipolar sum
Decomposing a plane wave into spherical harmonics, and then re-summing it, leads to a multipolar sum [75] which does not converge in the usual sense. The sum of our interest is the following
±$\mathcal{F}^{\pm}\sim\sum_{\ell=2}^{\infty}\Bigl{(}f^{\pm}_{\ell 2}Y^{(\pm 2)}_{\ell 2}+f^{\pm}_{\ell-2}Y^{(\pm 2)}_{\ell-2}\Bigr{)}\,,\qquad s=\pm 2\,.$ (39)
@@
Referring to the definition of the $f^{\pm}_{\ell m}$ in Eq. (28), notice how the terms with $m=s$ involve the difference between the contributions of the odd and even sectors, whereas the $m=-s$ involve the sum of both contributions. It should not come as a surprise, then, that while the first kind of terms converges rapidly in the usual sense, the second kind of terms does not.
@@
Figure 10: Plus (blue) and minus (red colors) contributions to the total amplification factor with positive spin weight $\mathcal{F}^{+}$, as a function of the highest angular harmonic included in the sum $\ell_{\rm max}$. Here we set $\theta_{\rm OL}=\pi/6,\phi_{\rm OL}=\pi/6,\omega M=0.5$. The contribution with $m=s$, in blue, converges rapidly in the usual sense. On the other hand, the contribution with $m=-s$, when summed in the usual way (purple dots), does not converge. However, by performing a Cesàro sum with order $\alpha=1$ (dashed, dark red line), or higher, e.g., $\alpha=2$ (solid red line), we can cast it into a convergent sum very accurately.
@@
For $\alpha=0$ this is just the usual sum, and for $\alpha=1$ this is the limit of the average of the partial sums. In Fig. 10 we show the behavior of both sequences of partial sums, as a function of the maximum value of $\ell$ included in the sum, for positive helicity (the same is observed for the negative helicity case). As expected, the $m=s$ converges rapidly in the usual sense. The $m=-s$ case does not converge in the usual sense (this is given by the order $\alpha=0$ Cesàro sum, shown as purple dots). By taking $\alpha=1,2$ (dashed and solid red lines), the sum becomes almost as rapidly convergent as its counterpart. Therefore we choose to separate the $m=\pm s$ contributions and sum one of them in the usual sense, and the other one in the Cesàro summation scheme with $\alpha$ sufficiently high to achieve the desired accuracy. In Fig. 11 we show the relative error in the calculation of $\mathcal{F}^{+}$ using two consecutive values of $\alpha$, defined as
@@
where the denominator is evaluated with $\alpha=5$ and $\ell_{\rm max}=60$. As clearly seen in the Fig. 11, the Cesàro summation is converging, and the error comitted is below $10\%$. More importantly, we know how increasing $\ell_{\rm max}$ diminishes the error committed. A more sophisticated numerical evaluation and analysis of the convergence of the resummation would be necessary to provide highly accurate lensed waveforms in the strong field regime. This is, nevertheless, beyond the scope of this work.
@@ Appendix B Relating distances in cosmology
While the form of these functions becomes more complicated for open or closed Universes (i.e. $\Omega_{k}\neq 0$), the Planck 2018 measurements of cosmological parameters were found to be consistent with a flat universe [55]. We therefore work under this assumption to simplify calculations of cosmological effects, and direct readers to other works to see the details of non-flat Universe models (c.f. [82, 83]).

### v3

@@ Appendix A Convergence of the multipolar sum
Decomposing a plane wave into spherical harmonics, and then resumming it, leads to a multipolar sum [75] that does not converge in the usual sense. The sum of our interest is the following
$\mathcal{F}^{\pm}\sim\sum_{\ell=2}^{\infty}\Bigl{(}f^{\pm}_{\ell 2}Y^{(\pm 2)}_{\ell 2}+f^{\pm}_{\ell-2}Y^{(\pm 2)}_{\ell-2}\Bigr{)}\,,\qquad s=\pm 2\,.$ (39)
@@
Referring to the definition of the $f^{\pm}_{\ell m}$ in Eq. (28), notice how the terms with $m=s$ involve the difference between the contributions of the odd and even sectors, whereas the m=sm=-s involve the sum of both contributions. It should not come as a surprise, then, that while the first kind of terms converges rapidly in the usual sense, the second kind of terms does not.
@@
Figure 10: Plus (blue) and minus (red colors) contributions to the total amplification factor with positive spin weight $\mathcal{F}^{+}$, as a function of the highest angular harmonic included in the sum $\ell_{\rm max}$. Here we set $\theta_{\rm OL}=\pi/6,\phi_{\rm OL}=\pi/6,\omega M=0.5$. The contribution with $m=s$, in blue, converges rapidly in the usual sense. On the other hand, the contribution with m=sm=-s, when summed in the usual way (purple dots), does not converge. However, by performing a Cesàro sum with order $\alpha=1$ (dashed, dark red line), or higher, e.g., $\alpha=2$ (solid red line), we can cast it into a convergent sum very accurately.
@@
For $\alpha=0$ this is just the usual sum, and for $\alpha=1$ this is the limit of the average of the partial sums. In Fig. 10 we show the behavior of both sequences of partial sums, as a function of the maximum value of $\ell$ included in the sum, for positive helicity (the same is observed for the negative helicity case). As expected, the $m=s$ converges rapidly in the usual sense. The m=sm=-s case does not converge in the usual sense (this is given by the order $\alpha=0$ Cesàro sum, shown as purple dots). By taking $\alpha=1,2$ (dashed and solid red lines), the sum becomes almost as rapidly convergent as its counterpart. Therefore we choose to separate the $m=\pm s$ contributions and sum one of them in the usual sense, and the other one in the Cesàro summation scheme with $\alpha$ sufficiently high to achieve the desired accuracy. In Fig. 11 we show the relative error in the calculation of $\mathcal{F}^{+}$ using two consecutive values of $\alpha$, defined as
@@
where the denominator is evaluated with $\alpha=5$ and $\ell_{\rm max}=60$. As clearly seen in the Fig. 11, the Cesàro summation is converging, and the error committed is below $10\%$. More importantly, we know how increasing $\ell_{\rm max}$ diminishes the error committed. A more sophisticated numerical evaluation and analysis of the convergence of the resummation would be necessary to provide highly accurate lensed waveforms in the strong field regime. This is, nevertheless, beyond the scope of this work.
@@ Appendix B Relating distances in cosmology
While the form of these functions becomes more complicated for open or closed universes (i.e., $\Omega_{k}\neq 0$), the Planck 2018 measurements of cosmological parameters were found to be consistent with a flat universe [55]. We therefore work under this assumption to simplify calculations of cosmological effects, and direct readers to other works to see the details of nonflat Universe models (cf. [82, 83]).